OpenStudy (anonymous):

please help through (2, 6) and perpendicular to y = -5/4x + 1. find equation

OpenStudy (anonymous):

slope will be 4/5 so eqn is, (y-y1)=slope(x-x1) (y-6)=4/5(x-2) 5(y-6)=4(x-2) 5y-30=4x-8 5y-4x=30-8 5y-4x=22

OpenStudy (anonymous):

slope will be 4/5 bcoz m1*m2=-1

OpenStudy (anonymous):

slope of first eqn is -5/4 so, (-5/4)*m2=-1 m2=4/5

OpenStudy (anonymous):

if u face any problem in understanding jst ask me.

OpenStudy (anonymous):

okay thanks